# Beyond Trend and Periodicity: Guiding Time Series Forecasting with Textual Cues

### v2

@@ 6 Experiments
We evaluate the TGForecaster model across four datasets to demonstrate the feasibility of the TGTSF task on the proposed datasets. As discussed in Section 5, each dataset is uniquely designed to test the task from different perspectives.
@@
TGForecaster is benchmarked against state-of-the-art (SOTA) methods including DLinear, FITS, PatchTST, iTransformer, and Time-LLM (Zeng et al., 2023; Xu et al., 2023; Nie et al., 2023; Liu et al., 2023; Jin et al., 2023). We specifically contrast it with the linear-based models DLinear and FITS to illustrate how TGForecaster surpasses periodicity-focused models. Comparisons with PatchTST highlight how textual information can enhance forecasting performance, even when using the same time series encoder. Time-LLM "reprograms" the LLaMa2 (Touvron et al., 2023), a large language model, which gives it capability of understanding text information. We use it as a baseline of multi-modal model. Specific settings differences, if any, are noted accordingly.
6.1 Evaluation on Toy Dataset
Experiment Settings
All of the models are following the same experimental setup with prediction length $H\in\{14,28,60,120\}$ and LBW length $T=60$.
@@ 6.1 Evaluation on Toy Dataset
Table 2 presents the performance comparison of various models on the toy dataset, with TGForecaster significantly outperforming all baseline models. Specifically, TGForecaster achieved an 80% improvement in Mean Squared Error (MSE) over the best performing transformer-based model, and a 96% increase over DLinear. These results underscore the considerable benefits of integrating textual guidance in TSF. Notably, a version of TGForecaster without news data, which lacks the ability to utilize auxiliary textual information, demonstrated substantially lower performance. This underscores the essential role of text in enhancing forecast accuracy.
Table 2: Forecasting result on toy and Electricity-Captioned dataset in MSE. The best result is highlighted in bold and the second best is highlighted in underline.
Dataset Pred. Len. TGForecaster
TGForecaster
w/o News
FITS DLinear PatchTST iTransformer TimeLLM
14 0.003 0.006 0.282 0.151 0.006 0.136 0.231
28 0.008 0.018 0.692 0.297 0.029 0.295 0.382
60 0.020 0.052 0.909 0.442 0.075 0.494 0.551
Toy 120 0.027 0.102 0.883 0.632 0.168 0.747 0.788
Pred. Len. TGForecaster
TGForecaster
lbw 120
FITS DLinear PatchTST iTransformer TimeLLM
96 0.124 0.127 0.134 0.140 0.130 0.148 0.131
192 0.144 0.146 0.149 0.153 0.149 0.162 0.152
336 0.160 0.164 0.165 0.169 0.166 0.178 0.160
Elec-c /Elec 720 0.193 0.200 0.203 0.204 0.210 0.225 0.192
Case Study
Figure 3 illustrates a segment of the toy dataset where the frequency changes within the forecasting horizon. In this visualization, the PatchTST model maintains the frequency observed in the look-back window, indicating an inability to adapt to new frequencies. Similarly, DLinear displays a collapsed pattern. Notably, without news input, TGForecaster shows the same behavior of PatchTST, suggesting that it relies solely on its time series encoder in the absence of textual cues. The models without external data also highlights the decreasing amplitude of predictions, a tendency to revert to safer, average predictions when faced with uncertainty in the far future—an issue known as the ’average shortcut’ phenomenon discussed in Section 2.1. Conversely, with the integration of external text information eliminating uncertainty, TGForecaster adeptly adapts to new frequencies at the appropriate moments, demonstrating the substantial benefits of incorporating textual data into the forecasting process.
Figure 3: Visualization on the toy dataset. Look-back window is 60. Orange line for prediction and blue line for groundtruth. This segment shows two frequency change points in the forecasting horizon. With the help of external text information, TGForecaster can accurately adapt to new frequency at the correct position, however other conventional time series models fail to do so.
6.2 Evaluation on Electricity-Captioned Dataset
Experiment Settings
We follow the experiment settings in previous works as follows: forecasting horizon $H\in\{96,192,336,720\}$, look back window length of 288. For fair comparison, we directly compare with the results report in the baseline original paper. And the TGForecaster is trained with the captioned version. On this dataset, we also test the impact of a shorter look-back window on the TGForecaster. We will report accordingly in Tab. 2.
@@ 6.2 Evaluation on Electricity-Captioned Dataset
As depicted in Tab. 2, TGForecaster demonstrates SOTA performance on the Electricity dataset, particularly effective at shorter forecasting horizons using minimal textual cues. However, TimeLLM slightly outperforms TGForecaster over longer forecasting periods. Given TimeLLM’s backbone is a large language model that also incorporates textual information, its edge may stem from its use of basic date and statistical data within its text inputs.
6.3 Evaluation on Weather-Captioned Dataset
Experiment Settings
We follow the experiment settings in previous works as follows: forecasting horizon $H\in\{96,192,336,720\}$, look back window length of 360. We trained all other models on the Weather-Medium and -large dataset with their setting on the original weather dataset accordingly. And the TGForecaster is trained with the captioned version. We will report accordingly in Tab. 3.
Table 3: Forecasting result on Weather-Midium and -Large dataset in MSE. The best result is highlighted in bold and the second best is highlighted in underscore.
Dataset Pred. Len. TGForecaster FITS DLinear PatchTST iTransformer TimeLLM
96 0.182 0.248 0.294 0.252 0.267 0.294
192 0.205 0.297 0.340 0.304 0.327 0.342
336 0.235 0.354 0.393 0.364 0.404 0.393
Weather -Medium 720 0.281 0.430 0.456 0.439 0.495 0.461
96 0.410 0.436 0.487 0.464 0.456 -
192 0.438 0.524 0.568 0.567 0.578 -
336 0.455 0.601 0.644 0.644 0.698 -
Weather -Large 720 0.497 0.692 0.725 0.745 0.832 -
Statistical Results
As demonstrated in Tab. 3, TGForecaster significantly outperforms other models across both the Weather-Medium and Weather-Large datasets, substantiating the efficacy of incorporating external text information in addressing the information insufficiency inherent in TSF models. Notably, despite its simplicity, the FITS model performs exceptionally well, second only to TGForecaster. This suggests that simply adding more training time series to complex models does not necessarily lead to improved performance, as more time series does not necessarily contain more information. The results highlight that strategic integration of textual data can provide a more substantial performance boost than merely increasing the quantity of time series data. Further, we also report the channel-wise performance in Appendix B. We notice a groundbreaking performance boost of over 60% on certain channels which were not predictable with historical time series data alone.
Table 4: Ablation result on Weather-Captioned-Medium in MSE.
@@ 6.3 Evaluation on Weather-Captioned Dataset
MiniLM
w/o Des.
MiniLM
w/o News

### v3

@@ 6 Experiments
FIATS is benchmarked against several state-of-the-art (SOTA) methods. These include linear-based models [7, 8] as , transformer-based models [1, 2], and fine-tuned LLM-based multimodal method [3]. Additionally, we compare pretrained time series "foundation models" [4, 6, 5]. This selection covers a range of approaches, including self-stimulated linear and nonlinear models, data-specific and pretrained models, LLM-based cross-modal models.
6.1 Evaluation on FM Toy Dataset
The FM Toy Dataset is generated using a fully-controlled frequency modulation system with a single intervention factor. This dataset has a theoretical error lower bound of 0, providing an ideal environment to analyze model performance.
@@ 6.1 Evaluation on FM Toy Dataset
Table 1 shows that FIATS achieves near-zero error, closely aligning with the theoretical lower bound, while other TSF methods, including pre-trained models, exhibit considerably higher error even with simple sinusoidal data. This confirms that the primary bottleneck is the self-stimulation assumption, not the quantity or variety of data.
Notably, linear-based models perform poorly in this intervention-sensitive context due to their limited parameters and linearity, resulting in collapsed predictions. In contrast, PatchTST, the best-performing self-stimulated model, demonstrates robustness in handling nonlinearity. As seen in Fig. 1, PatchTST captures periodicity, but with diminishing amplitude over time. This aligns with Proposition 2.1, which states that as the prediction horizon extends, the system’s behavior is more likely to be influenced by new interventions. As a result, self-stimulated models tend to predict conditional expectations in far future, leading to diminished amplitude. FIATS, by contrast, outperforms these models at longer horizons, as it incorporates the increasing influence of interventions while the impact of historical data fades.
Table 1: Forecasting result in MSE, comparing the intervention-aware FIATS against various TSF methods. The best result is highlighted in bold and the second best is highlighted in underscore.
Dataset
Pred.
Len.
FIATS FITS DLinear PatchTST iTrans. Chronos-L MOIRAI-L Time-MoE-U TimeLLM
14 0.003 0.282 0.151 0.006 0.136 0.012 0.013 0.012 0.231
28 0.008 0.692 0.297 0.029 0.295 0.047 0.062 0.035 0.382
60 0.020 0.909 0.442 0.075 0.494 0.129 0.133 0.107 0.551
FM Toy 120 0.027 0.883 0.632 0.168 0.747 0.374 0.385 0.295 0.788
96 0.124 0.134 0.140 0.130 0.148 0.154 0.152 0.149 0.131
192 0.144 0.149 0.153 0.149 0.162 0.177 0.171 0.168 0.152
336 0.158 0.165 0.169 0.166 0.178 0.197 0.192 0.183 0.160
Electricity Utility 720 0.190 0.203 0.204 0.210 0.225 0.242 0.236 0.229 0.192
96 0.182 0.248 0.294 0.252 0.267 0.293 0.299 0.258 0.294
192 0.205 0.297 0.340 0.304 0.327 0.357 0.356 0.318 0.342
336 0.235 0.354 0.393 0.364 0.404 0.448 0.457 0.413 0.393
Atmospheric Physics 2014-19 720 0.281 0.430 0.456 0.439 0.495 0.512 0.532 0.508 0.461
96 0.410 0.436 0.487 0.464 0.456 0.447 0.453 0.437 -
192 0.438 0.524 0.568 0.567 0.578 0.552 0.557 0.542 -
336 0.455 0.601 0.644 0.644 0.698 0.685 0.673 0.647 -
Atmospheric Physics 2014-24 720 0.497 0.692 0.725 0.745 0.832 0.754 0.765 0.734 -
6.2 Evaluation on Real World Dynamic System
@@
On the Electricity dataset, FIATS demonstrates SOTA performance, particularly effective at shorter forecasting horizons using minimal textual cues. It is interestingly to see that as the prediction length gets longer, the irregularly appeared holiday events tend to contribute less to the overall loss since the loss is averaged out. The TimeLLM with large-language-model backbone does show some capability to perform forecast according to such simple intervention information and obvious causal correlation.
Table 2: A selection of channel-wise performance on Atmos. Phy. 2014-19 dataset in MSE.
Channel FIATS FITS DLinear PatchTST iTrans. IMP.
p (mbar) 0.136 0.863 0.823 0.930 1.032 83.43%
Tpot (K) 0.182 0.316 0.352 0.322 0.353 42.18%
VPdef (mbar) 0.283 0.638 0.696 0.674 0.803 55.59%
rho (g/m³) 0.192 0.390 0.411 0.418 0.453 50.73%
raining (s) 0.790 0.873 0.937 0.859 0.994 8.04%
SWDR (W/m²) 0.182 0.308 0.385 0.296 0.377 38.39%
As shown in Table 1, FIATS consistently outperforms all baselines on the Atmospheric Physics dataset. These results underscore that incorporating external intervention information directly addresses the information insufficiency in conventional TSF models. In contrast, pretrained TSF models, such as Chronos-L and MOIRAI-L, despite being trained on larger datasets, still underperform FIATS—highlighting that scaling data alone cannot compensate for the fundamental limitations imposed by the self-stimulation assumption and the absence of intervention modeling.
Table 2 further breaks down the results by channel on the Atmospheric Physics dataset. FIATS shows substantial performance gains on variables such as pressure ($p$), air density ($\rho$), and vapor pressure (VPdef)—channels not directly referenced in the collected weather report. This demonstrates FIATS’s ability to perform channel-aware modeling and infer latent causal relationships between interventions and time series patterns, even when the correlation is indirect or not explicitly observed. The full breakdown performance shown in Appendix G.
6.3 Evaluation on Market System
We evaluate FIATS on the GAUD dataset to test its ability to handle real-world, intervention-driven market dynamics. Each time series tracks daily active users of a game over a 60-day input window and a 14-day forecast horizon. Due to large developer variability and temporal shifts, traditional TSF models struggle to generalize across games. FIATS addresses this by incorporating textual data, enabling intervention-aware forecasting.
Figure 3: Performance improvement with respect to the PatchTST on each time series in GAUD.
As shown in Fig. 3, FIATS consistently outperforms PatchTST, achieving an average improvement of 12.6% and ranking first on 59.6% of the games. In some cases, it boosts accuracy by up to 50–80%. Notably, for games released after 2021, where time series are short and cold-start issues arise, FIATS shows clear advantages. Its ability to generalize from textual interventions allows it to maintain performance where self-stimulated models like PatchTST fail to converge. Compared to TimeLLM, which depends on prompt templates, FIATS leverages raw textual semantics more effectively, resulting in broader applicability and better accuracy. Full results are provided in Appendix M.
6.4 Case Study & Ablation Study
Case Study: Visualization and Controllability Test
Fig. 4 visualizes three representative channels from the Atmospheric Physics dataset (full results in Appendix H). The first channel, atmospheric pressure ($p$), is sensitive to regional climate shifts but lacks strong short-term historical correlation. Its slow, subtle changes are challenging for traditional TSF models. PatchTST fails to capture these dynamics, defaulting to a flat prediction, while FIATS successfully models the trend by conditioning on relevant interventions.
Figure 4: Visualization of three channels on the 15,000th test sample of the Atmos. Phy. 2014-19 dataset. Blue indicates ground truth, Red shows FIATS, Green represents PatchTST, and Orange denotes FIATS with swapped interventions on the second and fourth forecast days. The CAPS causal alignment decoder exhibits distinct attention patterns across channels.
The second channel, rainfall duration (in seconds per 10 minutes), is sparse and lacks periodicity. PatchTST outputs near-zero values—its conditional expectation under uncertainty—while FIATS adjusts its predictions based on available intervention signals. It correctly forecasts the first rainfall event but misses the second due to misaligned or absent external information, reflecting dependence on accurate intervention input.
The third channel, solar radiation (SWDR), is not explicitly mentioned in the intervention but is indirectly influenced. FIATS captures its phase and amplitude accurately, thanks to the CASM design that enables cross-channel sensitivity modeling. PatchTST, by comparison, produces generic, misaligned waveforms.
A controllability test, Swapping interventions on the second and fourth forecast days confirms FlATS’s responsiveness. It updates predictions accordingly—forecasting rain on the fourth and clear skies on the second—demonstrating its ability to adapt to changes in external interventions in a causally aligned manner.
Case Study: Attention Map for Interpretability
: The CASM block analysis in Fig. 5 shows how the model focuses on different temporal features across layers. In the first layer, attention centers on the first sentence, providing temporal context for daily and annual periodicity. The second layer shifts attention to channel-specific signals, particularly the sixth sentence describing atmospheric pressure, reflecting the model’s sensitivity to channel-specific patterns and interventions. By the third layer, attention diversifies, focusing on relevant intervention aspects for each channel.
(a) CASM Layer1
(b) CASM Layer2
(c) CASM Layer3
Figure 5: Attention map of the CASM on the 15000th test sample of Atmos. Phy. 2014-19 dataset. We use three cross attention block in residual connection. The horizontal axis stands for channels and vertical stands for the 7 sentences of the weather report summary.
The CAPS causal alignment decoder, shown in Fig. 4, demonstrates distinct attention patterns across channels, highlighting the model’s ability to align time series data with textual interventions. Channels associated with periodic variables like SWDR exhibit clear periodicity in attention maps, indicating effective capture of cyclical patterns. Rainfall channel highlights historical rainfall, showcasing the model’s sensitivity to key moments. This adaptability is driven by CASM, enabling the model to tailor its attention based on each channel’s unique characteristics. Analysis on such attention map may further reveal the underlying causal correlation about how a time series or a system reacts to certain external intervention for future work. Full analysis see Appendix K.
Table 3: Ablation result on Atmos. Phy. 2014-19 in MSE.
@@ 6.4 Case Study & Ablation Study
Zero
Desc.
Zero
News